OpenStudy (anonymous):

a man is 4 times as old as his son. after 16 years he will be only twice as old as his son. Find their present ages

terenzreignz (terenzreignz):

I don't know what this is doing in differential equations, but the first thing to do in this type of problem is to 'represent'. So, we can let x be the current age of the son. If that's the case, what's the age of the father?

OpenStudy (rituraj):

Let age of son = x So age of father = 4x a/q: 4x + 16 = 2 (x + 16) So x = 8 Age of son = 8 Age f father = 32